# Optimal and fast online change point estimation in linear regression

### v1

@@ 2.1 Risk bounds
The idea of proof (for details see Section 5) is to characterize an event on which the detector reports a change point without false alarms and achieves a detection delay that is in the range of the considered rate; then such an event is shown to occur with a high probability. Theorem 1 provides an upper bound on the detection risk for online detectors in our model. The constants are explicit, but they are not optimized and could potentially be improved. As demonstrated later in Section 3, the rates in Theorem 1 are also achieved by lower bounds, revealing that FLOC is minimax optimal in detection of the change point.
@@ 2.2 Computational complexity
A pseudocode implementation of FLOC is provided in Algorithm 1. The algorithm maintains sums and weighted sums in bins of fixed size to calculate the test statistics. When the third bin is full, the first bin is discarded, and a new bin is initiated. The benefits of varying window sizes over a fixed window size become evident in the implementation of FLOC. While a constant window size achieves the same asymptotic properties, it would require storing individual observations to allow for gradual updates as the window moves. As the window size depends on $n$, the storage cost for a fixed window size would grow with the rate of observations. In contrast, Algorithm 1 achieves computational and memory costs of $\mathcal{O}(1)$, independent of the window size or the number of observation until the current time. This property is particularly advantageous for online detectors, as minimal computational and memory requirements are often considered essential design criteria (cf. Introduction).
@@
For the detector FLOC as defined in (3b)–(3f), the following holds:
@@
    The computational cost is $\mathcal{O}(1)$.
@@
FLOC simultaneously monitors for both jumps and kinks by leveraging the stored sums in the computation of test statistics for both types of change. To focus on detecting only one type, we can formally set the threshold for the other to infinity. A practical challenge in Algorithm 1 is the choice of the bin sizes $N_{J},N_{K}$ and the thresholds $\rho_{J},\rho_{K}$, as these parameters would influence the empirical performance. The parameter values used in the proof of Theorem 1 can serve as a guideline, though they are mainly tuned for their asymptotic performance. We discuss the choice of parameters in Section 4.1. The implementation can also be adapted to incorporate a known pre-change signal $f_{-}$ by directly setting this in place of its estimate $\hat{f}_{-}$.
@@
5:compute scaling factor $d$ $\leftarrow\frac{(2N_{K}+r_{K}+1)(2N_{K}+r_{K}+2)(4N_{K}+2r_{K}+3)}{6}$
@@ 3 Lower risk bounds
In this section we complement the upper bounds in the previous section with matching lower bounds, and discuss the relation to existing results in the literature.
@@
Assume the segmented linear regression model (1a)–(1c). Then, there exist positive constants $r_{*J}$ and $r_{*K}$, independent of $n$, such that:
@@
and
$\liminf_{n\to\infty}\inf_{\hat{\tau}_{n}\in\mathcal{T}}\max_{\theta=(\tau,\alpha_{-},\alpha_{+},\beta_{-},\beta_{+})\in\Theta^{K}_{\delta_{0}}}\mathbb{E}_{\tau}\left[\left(\frac{n^{1/3}}{\log^{1/3}n}\left(\hat{\tau}_{n}-\tau\right)\right)^{2}\right]\geq r_{*K}>0.$
@@
In combination with Theorem 1, Theorem 3 shows that the two distinct rates for online detection of jumps and kinks are minimax optimal. These rates closely align with the offline setup, differing only in log factors (see Table 1 in the Introduction). Notably, FLOC attains the minimax optimality in detection of both jumps and kinks. For the detection of a jump, the optimal rate for the segmented linear regression model coincides with the optimal rate for the submodel with piecewise constant signal and a single jump (see e.g. Chapter 6 in Korostelev and Korosteleva,, 2011). It reveals that asymptotically a change in the slope is negligible in comparison to a jump. Further, for the submodel with piecewise constant signal, Yu et al., (2023) showed a rate of order $\log\frac{n}{\alpha}$ on the detection delay under the condition $\mathbb{P}\left(\hat{\tau}<\infty\mid\text{no change}\right)\leq\alpha$, which ensures that under the null hypothesis the false alarm probability of online detector is at most $\alpha$. The computational cost for their proposed detector scales linearly with the number of observations available at time $t$, and this complexity was later improved to $\mathcal{O}(\log t)$ using functional pruning (Romano et al.,, 2023). In contrast, our FLOC detector, designed for a more general model, attains the same detection delay rate in this specific submodel, and requires a computational and memory costs independent of time $t$, see also Section 1.2.
@@ 4 Numerical experiments
In this section, we first introduce an empirical method for the selection of thresholds in FLOC (Algorithm 1). Then we evaluate the empirical performance of FLOC on simulated data in regard to the expected detection delay and false alarm probability. In addition, we apply FLOC to analyze the Covid-19 excess death data from the United States. The codes we use in this section are available as an R package at GitHub under https://github.com/AnnikaHueselitz/FLOC.
@@
Figure 3: Expected detection delay for a true change point at observation $\tau_{0}=500$ for bin sizes 5 (blue solid curve) and 10 (red dashed curve) for a jump (left) and a kink (right). The thresholds were tuned such that each achieves a false alarm probability of $\mathbb{P}\left(\hat{\tau}\leq 1000\right)\approx 0.5$. Here $f_{-}$ was estimated from $k=500$ observations, for the threshold tuning $r=1000$ sets of data were generated and the estimation of the expected detection delay was done with $500$ repetitions.
@@
Table 2: Performance of FLOC (Algorithm 1) on simulated data. The thresholds $\rho_{J}$ and $\rho_{K}$ are tuned such that a false alarm probability (FA) of $0.5$ is reached with $k$ historical data points using 10000 repetitions. The false alarm probability and the expected detection delay, with a change point at $\tau=k$ for jump sizes 2, 1 and 0.5 and for changes in the slope of sizes 0.5, 0.1 and 0.02 are estimated through Monte Carlo simulations with 200 repetitions.

### v2

@@ 2.1 Risk bounds
The idea of the proof (for details see Section 5) is to characterize an event on which the detector reports a change point without false alarms and achieves a detection delay that is in the range of the considered rate; then such an event is shown to occur with a high probability. The constants in Theorem 1 are explicit, but they are not optimized and could potentially be improved.
The originally assumed requirement of $k\asymp n$ historical data points can be substantially relaxed. As shown in the proof in Section 5, it suffices to choose $k\asymp n^{2/3}\log(n)^{1/3}$ in the jump case, and $k\asymp n^{8/9}\log(n)^{1/9}$ in the kink case. In the special case of a piecewise constant model, i.e., when $\beta_{-}=\beta_{+}=0$ in (1b), the requirement can be further reduced to $k\asymp\log(n)$. The gap between the requirements $k\asymp n^{2/3}\log(n)^{1/3}$ and $k\asymp\log(n)$ reveals the additional sample complexity of historical data incurred in the more general piecewise linear setting.
This requirement on $k$ has practical implications when applying the online detector to monitor multiple change points. In online settings, the method can be naturally extended to the multiple change point scenario by restarting the detector each time a change is declared. In such a framework, the minimum segment length between two change points must be at least of order $k$.
As demonstrated later in Section 3, the rates established in Theorem 1 are matched by corresponding lower bounds, revealing that $\hat{\tau}_{J,n}$ is minimax optimal for detecting a jump, while $\hat{\tau}_{K,n}$ is minimax optimal for detecting a kink. Further, FLOC, defined as the minimum of these two, is minimax optimal in detection of the change point, simultaneously for jump and kink types of structural changes. Motivated by this observation, one might naturally consider inferring the type of structural change, either a jump or a kink, by examining which of the two detectors $\hat{\tau}_{J,n}$ and $\hat{\tau}_{K,n}$ triggers an alarm first. The following proposition establishes that this approach is indeed theoretically justified.
Proposition 2.
Assume the same setup as in Theorem 1 and let $n$ be sufficiently large. Then:
$\mathbb{P}_{\tau}\left(\hat{\tau}_{K,n}<\hat{\tau}_{J,n}\right)\leq 2n^{-3}\qquad\text{if}\quad\theta=(\tau,\alpha_{-},\alpha_{+},\beta_{-},\beta_{+})\in\Theta^{J}_{\delta_{0}},$
and $\mathbb{P}_{\tau}\left(\hat{\tau}_{J,n}<\hat{\tau}_{K,n}\right)\leq 2n^{-3}\qquad\text{if}\quad\theta=(\tau,\alpha_{-},\alpha_{+},\beta_{-},\beta_{+})\in\Theta^{K}_{\delta_{0}}.$
The proof (see Section 5) builds on the high probability events introduced in the proof of Theorem 1, under which the false alarm rates of both detectors are well controlled. In the presence of a jump, the kink detector reacts more slowly due to its larger bin size. Conversely, when a kink is present, the induced change in slope is too gradual to trigger the jump detector before the kink detector detects the change. The rate of upper bound $n^{-3}$ could be replaced by $n^{-C}$ for any constant $C\geq 0$, while doing so would require enlarging the bin sizes, which would in turn worsen the constants $r_{J}^{*}$ and $r^{*}_{K}$ in Theorem 1. This result confirms that each detector is very unlikely to falsely signal first in the incorrect regime.
Furthermore, using the decomposition in (2), Theorem 1 yields statistical guarantees on the false alarm probability, the average run length and the expected detection delay (cf. Section 1.2).
Corollary 3.
Assume the same setup as in Theorem 1 and let $n$ be sufficiently large. Then:
  1. i.
    For the jump detector $\hat{\tau}_{J,n}$, it holds that $\mathbb{P}_{\tau}(\hat{\tau}_{J,n}<\tau)\leq n^{-3}$, $\mathbb{E}_{0}\left[\hat{\tau}_{J,n}\right]\geq 1-n^{-3}$ and
    $\sup_{\theta=(\tau,\alpha_{-},\alpha_{+},\beta_{-},\beta_{+})\in\Theta^{J}_{\delta_{0}}}\frac{n}{\log n}\mathbb{E}_{\tau}\left[\hat{\tau}_{J,n}-\tau\mid\hat{\tau}_{J,n}\geq\tau\right]\;\leq\;\sqrt{r^{*}_{J}}<\infty.$ (6a)
  2. ii.
    For the kink detector $\hat{\tau}_{K,n}$, it holds that $\mathbb{P}_{\tau}(\hat{\tau}_{K,n}<\tau)\leq n^{-4}$, $\mathbb{E}_{0}\left[\hat{\tau}_{K,n}\right]\geq 1-n^{-4}$ and
    $\sup_{\theta=(\tau,\alpha_{-},\alpha_{+},\beta_{-},\beta_{+})\in\Theta^{K}_{\delta_{0}}}\frac{n^{1/3}}{\log^{1/3}n}\mathbb{E}_{\tau}\left[\hat{\tau}_{K,n}-\tau\mid\hat{\tau}_{K,n}\geq\tau\right]\;\leq\;\sqrt{r^{*}_{K}}<\infty.$ (6b)
  3. iii.
    For the FLOC detector $\hat{\tau}_{n}$, we have $\mathbb{P}_{\tau}(\hat{\tau}_{n}<\tau)\leq n^{-3}+n^{-4}$ and $\mathbb{E}_{0}\left[\hat{\tau}_{n}\right]\geq 1-n^{-3}-n^{-4}$, and the same upper bounds as in (6a) and (6b) hold.
As noted for Proposition 2, the bounds of $n^{-3}$ and $n^{-4}$ may be replaced by $n^{-C}$ for any constant $C\geq 0$, provided appropriate parameter values are chosen. In this regard, the parameters can be tuned to achieve a user-specified false alarm probability or average run length, see Section 4.1 for details.
In addition, it is worth noting that, in Theorem 1, Proposition 2 and Corollary 3, the Gaussian assumption on the noise is used solely to control the deviation of weighted sums of the noise variables. More precisely, in the proofs in Section 5, we employ the concentration inequality
$\mathbb{P}\left(\left|\sum_{i=1}^{n}w_{i}\varepsilon_{i}\right|\geq t\right)\leq\exp\left(-\frac{t^{2}}{2}\right)\qquad\text{for}\quad t>0,$
where the weights $w_{i}$ are deterministic and satisfy $\sum_{i=1}^{n}w_{i}^{2}=1$. Similar concentration bounds hold for sums of independent sub-Weibull random variables, see e.g. Kuchibhotla and Chakrabortty, (2022). Further, certain forms of temporal dependence can be incorporated using the functional dependence framework (introduced in Wu,, 2005). As a consequence, the established statistical guarantees of our proposed online detectors can be extended to these more general noise settings.
@@ 2.2 Computational complexity
A pseudocode implementation of FLOC is provided in Algorithm 1. The algorithm maintains sums and weighted sums in bins of a fixed size to calculate the test statistics. When the third bin is full, the first bin is discarded, and a new bin is initiated. The benefits of varying window sizes over a fixed window size become evident in the implementation of FLOC. While a constant window size achieves the same asymptotic properties, it would require storing individual observations to allow for gradual updates as the window moves. As the window size depends on $n$, the storage cost for a fixed window size would grow with the rate of observations. In contrast, Algorithm 1 achieves computational and memory costs of $\mathcal{O}(1)$, independent of the window size and the number of observation until the current time. This property is particularly advantageous for online detectors, as minimal computational and memory requirements are often considered essential design criteria (cf. Introduction).
@@
For the FLOC detector as defined in (5b)–(5f), the following holds:
@@
    The computational cost is $\mathcal{O}(1)$ per incoming observation.
@@
We note that a full batch computation of $\hat{f}_{-}$ in (5a) using the historical data of size $k$ requires $\mathcal{O}(k)$ runtime, but this is performed only once. Alternatively, $\hat{f}_{-}$ can be updated sequentially, resulting in an $\mathcal{O}(1)$ computational cost per data point.
@@
5:compute scaling factor $d$ $\leftarrow{(2N_{K}+r_{K}+1)(2N_{K}+r_{K}+2)(4N_{K}+2r_{K}+3)}/{6}$
@@ 3 Lower risk bounds
In this section, we complement the upper bounds in the previous section with matching lower bounds, and discuss the relation to existing results in the literature.
@@
Assume the segmented linear regression model (1a)–(1c). Then, there exist positive constants $r_{*J}$ and $r_{*K}$, independent of $n$, such that
@@
and $\liminf_{n\to\infty}\inf_{\hat{\tau}_{n}\in\mathcal{T}}\max_{\theta=(\tau,\alpha_{-},\alpha_{+},\beta_{-},\beta_{+})\in\Theta^{K}_{\delta_{0}}}\mathbb{E}_{\tau}\left[\left(\frac{n^{1/3}}{\log^{1/3}n}\left(\hat{\tau}_{n}-\tau\right)\right)^{2}\right]\geq r_{*K}>0.$
@@
In combination with Theorem 1, Theorem 5 shows that the two distinct rates for online detection of jumps and kinks are minimax optimal. These rates closely align with the offline setup, differing only in log factors (see Table 1 in the Introduction). Notably, FLOC attains the minimax optimality in detection of both jumps and kinks. For the detection of a jump, the optimal rate for the segmented linear regression model coincides with the optimal rate for the submodel with piecewise constant signal and a single jump (see e.g. Chapter 6 in Korostelev and Korosteleva,, 2011). It reveals that asymptotically a change in the slope is negligible in comparison to a jump. Further, for the submodel with piecewise constant signal, Yu et al., (2023) showed a rate of order $\log\frac{n}{\alpha}$ on the detection delay under the condition $\mathbb{P}\left(\hat{\tau}<\infty\mid\text{no change}\right)\leq\alpha$, which ensures that under the null hypothesis the false alarm probability of online detector is at most $\alpha$. The computational cost for their proposed detector scales linearly with the number of observations available at time $t$, and this complexity was later improved to $\mathcal{O}(\log t)$ using functional pruning (Romano et al.,, 2023). In contrast, our FLOC detector, designed for a more general model, attains the same detection delay rate in this specific submodel, and requires a computational and memory costs independent of time $t$, see also Section 1.2 and Corollary 3.
@@ 4 Numerical experiments
In this section, we first introduce an empirical method for the selection of thresholds in FLOC (Algorithm 1). Subsequently, we assess the empirical performance of FLOC on simulated datasets by benchmarking it against state-of-the-art online change point detection methods. In addition, we investigate the robustness of FLOC in settings that go beyond the scope of our theoretical guarantees, particularly in the presence of non-Gaussian noise. Finally, we demonstrate the application of FLOC to the analysis of COVID-19 excess mortality data from the United States. The implementation used in this section is publicly available as an R package on GitHub at https://github.com/AnnikaHueselitz/FLOC.
4.1 Choice of parameters
FLOC involves five tuning parameters. Theorem 1 provides theoretical guidance for their selection: the historical data size $k=cn$, with $c<\delta_{0}$ (i.e., smaller than the minimal detectable change), thresholds $\rho_{J}=4c/5$ and $\rho_{K}=4c/(5n)$ and bin sizes $N_{J}=10^{3}\log(n)/(2c^{2})$ and $N_{K}=(300/c^{2})^{1/3}n^{2/3}\log^{1/3}(n)$. However, these settings may be impractical in cases where $\delta_{0}$ is unknown. We therefore provide empirical recommendations below, based on simulation studies.
The amount of historical data $k$ governs the accuracy of estimating the pre-change distribution. Larger values of $k$ yield more reliable estimates and result in a more stable detector. See Table 2 for simulation results illustrating the effect of $k$ on detection delay. In practice, one should utilize as much historical data as is available or computationally feasible. When $k$ is small, higher thresholds are needed to control the type I error, which in turn increases the detection delay. We can also incorporate a known pre-change signal $f_{-}$ by directly substituting it in place of its estimate $\hat{f}_{-}$.
In general, increasing the detection threshold reduces the typeI error at the cost of increasing the typeII error. The type I error is typically measured by the false alarm probability $\mathbb{P}_{\tau}\left(\hat{\tau}<\tau\right)$, which is the probability that a false alarm, when the change point is at $\tau$, or by the average run length under the null hypothesis $\mathbb{E}_{0}\left[\hat{\tau}\right$], which is the expected run length if no change occurs. The type II error is often measured by the expected detection delay $\mathbb{E}_{\tau}\left[\hat{\tau}-\tau\mid\hat{\tau}\geq\tau\right]$. In applications, $n$ does not need to be specified beforehand. Thus, we will use in the following the expected detection delay and the average run length in terms of the number of observations.
@@ 4.1 Choice of parameters
Figure 3: Comparison of the type I error, expressed as the false alarm probability $\mathbb{P}(n\hat{\tau}_{n}<1000)$, and the type II error, measured by the expected detection delay for a change point at observation $500$, for the jump detection (left) and the kink detection (right) using FLOC. In both plots, the bin size is $N_{J}=N_{K}=5$ and $f_{-}$ is estimated using $k=500$ observations. Threshold tuning is performed with $r=5000$ data sets and the expected detection delay is estimated with $200$ repetitions. The jump size used is $1$ and the change in slope is $0.01$.
Figure 3 illustrates the relationship between the false alarm probability and the expected detection delay. In practice, thresholds can be tuned to control the type I error at a user-specified level $\eta$ e.g. by bounding the false alarm probability. This amounts to finding the smallest threshold such that $\mathbb{P}_{\tau}\left(\hat{\tau}<\tau\right)\leq\eta$ for some fixed $\tau$. More precisely, for given bin sizes $N_{J}$ and $N_{K}$ and $k$ historical observations, we generate independently $r$ data sets $X_{1}^{(m)},\ldots,X_{k+n\tau}^{(m)}$ for $1\leq m\leq r$ under the null hypothesis. For each data set, we compute the maximum statistics $X^{*}_{J,m}\coloneqq\max_{1\leq i<n\tau}J_{i}^{(m)}$ for jump detection and $X^{*}_{K,m}\coloneqq\max_{1\leq i<n\tau}K_{i}^{(m)}$ for kink detection. By default, FLOC simultaneously monitors for both a jump and a kink; however, it can be configured to detect only one of these changes by setting the threshold $\rho_{K}$ (or $\rho_{J}$) to infinity, thereby disabling the detection of kinks (or jumps). When using FLOC to detect only jumps , we select the jump threshold $\rho_{J}$ as the $\left(1-\eta\right)$-quantile of $\{X^{*}_{J,m}:1\leq m\leq r\}$. A similar procedure applies when detecting only kinks. When FLOC is applied to monitor both change types simultaneously, the thresholds can be selected as appropriate quantiles of $\{X^{*}_{J,m}:1\leq m\leq r\}$ and $\{X^{*}_{K,m}:1\leq m\leq r\}$ such that $\mathbb{P}_{\tau}\left(\hat{\tau}_{J}<\tau\right)\approx\mathbb{P}_{\tau}\left(\hat{\tau}_{k}<\tau\right)$ and $\mathbb{P}_{\tau}\left(\hat{\tau}_{J}<\tau,\hat{\tau}_{k}<\tau\right)\approx\eta$. The effectiveness of this approach is examined by simulation in Table 2. A related strategy can be employed to calibrate thresholds with respect to the average run length. Due to the rolling window structure, FLOC exhibits approximate memorylessness under the null, suggesting that the run length distribution is roughly exponential. To target an average run length of at least $\mathbb{E}_{0}\left[\hat{\tau}\right]\geq\tau$, one can use the same threshold selection procedure as for the false alarm probability, with $\eta=1-1/e$. Table 3 reports thresholds obtained via this approach along with the empirical average run lengths that are actually achieved.
@@
Table 2: Performance of FLOC (Algorithm 1) on simulated data. The thresholds $\rho_{J}$ and $\rho_{K}$ are calibrated with $10000$ repetitions to achieve a target false alarm probability (FA) of $0.5$, when the change point is at observation $1000$ for $k\in\{500,1000\}$ historical observations, and the change point at $10000$ for $k=5000$. The actual false alarm probability and the expected detection delay in terms of number of observations (last six columns), under a change point at $k$-th observation, are estimated via Monte Carlo simulations with 200 repetitions, across varying magnitudes of jumps and kinks.